a horizontal force of 150N is used to push a 40kg packing crate a distance of 6m on a rough horizontal surface. if the crate moves at constant speed, find the workdone by the 150N force and the coefficient of kinetic friction between the crate and surface
Answer #1

We know, work done = F x , where , F= force applied, x= distance moved.
now, F= 150 N, x= 6 m.
So , work done = 150 x 6 = 900 J

Now, as there is no acceleration , so , force applied=force of kinetic friction.
So, 150 N = force of kinetic friction = μmg = μ x 40 x 9.8

So, μ =150/(40 x 9.8) = 0.38

answered by: S.O.S

> (150cos15 )/ (40 x 9.8 - 150 sin 15) = μ

Egan Beach Tue, Feb 1, 2022 7:13 PM